### v1

# DeepUniUSTransformer: Towards A Universal UltraSound Model with Prompted Guidance

Abstract
Ultrasound is a widely used imaging modality in clinical practice due to its low cost, portability, and safety. Current research in general AI for healthcare focuses on large language models and general segmentation models, with insufficient attention to solutions addressing both disease prediction and tissue segmentation. In this study, we propose a novel universal framework for ultrasound, namely DeepUniUSTransformer, which is a promptable model accommodating multiple clinical task. The universality of this model is derived from its versatility across various aspects. It proficiently manages any ultrasound nature, any anatomical position, any input type and excelling not only in segmentation tasks but also in computer-aided diagnosis tasks. We introduce a novel module that incorporates this information as a prompt and seamlessly embedding it within the model’s learning process. To train and validate our proposed model, we curated a comprehensive ultrasound dataset from publicly accessible sources, encompassing up to 7 distinct anatomical positions with over 9.7K annotations. Experimental results demonstrate that our model surpasses both a model trained on a single dataset and an ablated version of the network lacking prompt guidance. We will continuously expand the dataset and optimize the task specific prompting mechanism towards the universality in medical ultrasound. Model weights, datasets, and code will be open source to the public.
Keywords:
Promptable Learning Ultrasound Universal Model.
1 Introduction
Ultrasound imaging plays an important role in clinical applications, and is known for its economy, safety and portability. In the past, specialist ultrasound image models based on deep learning have made significant progress, but their capability scenario are usually narrow and are limited by the amount of data annotation. Recently, General Medical Artificial Intelligence (GMAI) models [15, 21, 18, 12] have gained momentum. Unlike their predecessors relying on single dataset or task, these models demonstrate notable performance in generalizing to a wide range of tasks and data. However, a universal model for tailored for ultrasound modalities have not undergone thorough exploration yet. This necessitates a thorough evaluation of ultrasound images, taking into account the nature of the image (tumor or non-tumor), different positions, diverse task and varied input type.
To train a universal model applicable to the ultrasound, we have curated a collection of publicly available datasets from online sources. This dataset spans 7 different anatomical positions, including breast, thyroid, appendix, kidneys, heads, cardiac, livers, as illustrated in Figure 1. Termed as BroadUS-9.7K, the dataset comprises over 6.9K ultrasound images and more than 9.7K annotations, including classifications and segmentations. We intend to release this dataset to serve as a benchmark for the medical ultrasound GMAI community, fostering the development and evaluation of universal models. Designing a versatile model capable of handling such a diverse dataset poses a novel challenge. The incorporation of suitable prompt guidance holds the potential to enhance the model’s adaptability, enabling it to address a spectrum of tasks.
Figure 1: The BroadUS-9.7K dataset contains 9.7K annotations of 6.9K ultrasound images from 7 different anatomical positions. (a) The number of effective instances corresponding to nature of the image, position, task and input type. Note that a breast image can contain both segmentation and classification labels, and an image with segmentation labels can form three different input types (b) The different anatomical positions, and their corresponding public dataset abbreviations.
In recent years, significant strides have been made in the development of AI models for ultrasound imaging, addressing diverse tasks such as breast tumor prediction [10], thyroid tumor segmentation [13], COVID-19 prognosis [19], and prenatal image quality control [7]. However, these specialized models face challenges when confronted with new tasks or datasets, requiring considerable effort for retraining from scratch. This approach neglects the intrinsic relationships between datasets and tasks. Currently, the focus is shifting towards universal models that aim to uncover these inherent connections. Zhang et al. [23] proposed leveraging CLIP embeddings to guide the segmentation of multi-organ models. Han et al. [6] introduced the concept of target code, incorporating prompts into multi-sequence MRI generation models through HyperConv layer. Nevertheless, these models exhibit certain shortcomings. Firstly, the incorporation of prompts directly into the mid-feature layer of the encoder-decoder lacks direct guidance for the network’s intrinsic learning process. Secondly, incorporating fixed one-hot encoding and CLIP features without learnable parameters may makes network optimization challenging.
In this study, we propose a novel universal framework for ultrasound, termed as DeepUniUSTransformer, which is a promptable model accommodating multiple clinical tasks. Our contributions are threefold. Firstly, we have collected a rich set of publicly available ultrasound datasets containing a variety of anatomical positions, on which validation of our model is convincing. Secondly, we introduce a universal model framework built upon the transformer architecture. This framework, rooted in four different model prompts (i.e. nature of the image, positions, task and input type), significantly enhances the network’s adaptability across diverse circumstances. The model can adapts to new tasks and categories by adjusting the prompt and utilizing the existing structure without the need for additional branches. Thirdly, we demonstrate the effectiveness and universality of our framework by conducting extensive experiments on the BroadUS-9.7K dataset. Our model surpasses both a model trained on a single dataset and an ablated version of the network lacking prompt guidance. Additionally, we conduct a zero-shot experiment to showcase the model’s capacity for generalization to novel unseen dataset.
2 Methodology
Figure 2: The architecture of the proposed DeepUniUSTransformer.
The overall architecture of DeepUniUSTransformer is shown in Fig. 2. We propose a general model that uses prompts to perform multiple tasks including segmentation and classification on ultrasound images simultaneously. Our model is based on an encoder-decoder structure. The encoder is responsible for extracting discriminating features, and different decoders are responsible for different tasks which are augmented by prompt. Four kinds of prompt (as shown in the Fig. 2), which respectively prompt the nature, position, task and type of the ultrasound image, will be added to each transformer layer by means of prompt projection embedding to enhance the versatility and performance of the network.
2.1 Uni-Transformer Learning Framework
The proposed DeepUniUSTransformer learning framework, which is designed to perform multiple tasks on ultrasound images simultaneously. The primary part of the network is mainly modified based on Swin-Unet [3]. It comprises one encoder and two decoders. The encoder is responsible for extracting features, and the two decoders are responsible for segmentation and classification tasks respectively. Note that the decoder branch responsible for segmentation incorporates a skip connection originating from the encoder (same as the swin-unet). Compared with the decoder branch responsible for classification and the decoder branch responsible for segmentation, the upper sampling layer is deleted, but the number of layers is kept consistent, so as to ensure computational equivalence among different task branches.
Inspired by [24], which are also trained on multiple publicly available datasets. Since the amount of data in each dataset was inconsistent, we carried out data balancing based on position. This strategy aims to optimize the model’s computational learning for distinct domain knowledge. Furthermore, following the principles of curriculum learning, we prioritizes segmentation data, succeeded by classification data in each training round. This sequential learning approach capitalizes on the object’s edge and location information acquired through segmentation, may enhancing subsequent image classification tasks. Regarding training losses, we employ two distinct types for segmentation tasks: the dice loss and cross-entropy loss. Specifically, the classification task is optimized using cross-entropy loss.
2.2 Prompting Strategy
We employed four distinct prompts to provide the model with diverse information regarding images and tasks: nature, position, task, and type prompts. The nature prompt serves as a binary indication, discerning whether the current image is an organ anatomy image or a tumor image. Given the substantial disparity between the two, anatomical images prioritize structural regularity, whereas tumor images demand attention to information heterogeneity. The position prompt designates the anatomical location to which the image belongs, offering valuable insights related to domain knowledge. The task prompt indicates the current task assigned to the network (segmentation, classification), providing clarity regarding the network’s preference for extracting image features. Lastly, the type prompt specifies the image’s input type, encompassing whole, local, and location. Introducing this prompt facilitates the adjustment of spatial granularity in image pattern recognition and maximizes the utilization of location information to aid the learning. For this, we utilize data with segmentation labels, generating additional crop (local type) and local brightness enhancement input (location type).
We drew inspiration from the position embedding technique used in Vision Transformers (ViT) [4], and our implementation strategy for prompts is both simple and intuitive. Firstly, we define four corresponding prompt one-hot vectors based on the input image. These vectors are then concatenated and projected into the same dimensions as the the transformer layer using a fully connected (fc) layer. Subsequently, the projected prompt vectors are directly added to the feature $x$ that would have been processed by the successive transformer layer. It is important to note that we perform this operation at each layer of the decoder transformer to ensure that the prompts are fully involved in the task resolution process. The projection parameters of the fc layer are not shared across layers. Furthermore, since the fc layer is learnable, it automatically adapts and optimizes the projection method based on different feature learning process. This design is remarkably simple and significantly enhances the reusability of the network structure. When encountering new task datasets in different domains, we only need to modify the information in the prompts without adding new branches or task head. Specifically, the implementation can be expressed as follows:
$x^{{}^{\prime}}=x+fc_{i,k}(concat(p_{nature},p_{position},p_{task},p_{type})),i=1,2,3,k=1,2$ (1)
whereas $x_{i,k}$ is the input feature, $p$ is the prompt vector, $fc$ is the fully connected layer, $i$ and $k$ represent the $i$-th layer and the $k$-th branch (segmentation and classification) respectively.
2.3 Shifted Domain Testing and Fine-tuning
To explore the ability of proposed universal model on new domain dataset, we also propose a LoRA fine-tuning scheme inspiration from LoRA [9] in the natural language research domain. We propose freezing all parameters of the entire encoder-decoder structure and only fine-tuning the learnable fully connected projection layer parameters when prompts are introduced to adjust the universal model for new task/domain.
3 Experimental Results
Dataset and Implementation Details. The details of the dataset can be seen in Table 1. In some publicly available datasets, there are cases where a patient has multiple images. When partitioning the data, we ensure that all images of the same patient are included in the same partition. The dataset was divided into training, validation, and testing sets in a ratio of 7:1:2. The segmentation loss consists of a 0.4 ratio of cross-entropy loss and a 0.6 ratio of dice loss. We utilized the AdamW optimizer with an initial learning rate of 3e-4 and trained the models for 200 epochs in each experiment. Data augmentation techniques such as random horizontal flipping, random rotation within the range of -20° to 20°, and random cropping were applied. The code implementation was carried out using PyTorch, with the NVIDIA A4000 GPU utilized for computational acceleration. In this paper, the evaluation metrics for segmentation employed the dice coefficient, while the classification evaluation employed accuracy. Both metrics range from 0 to 1, with higher values indicating better performance.
Quantitative Analysis. We conducted two main experiments in our study. The first experiment aimed to compare the results of training a single dataset alone with the results obtained from a multi-task framework without prompts, as well as using our proposed prompt method. The second experiment focused on exploring the generalization performance of our model. For this purpose, the BUSI dataset was kept separate and not used in the fusion and prompt experiments. Specifically, we compared the performance of a single model trained solely on the BUSI dataset (labeled as Scratch in Table 3), the performance of a single model from BUS-BRA (another breast tumor dataset) for zero-shot testing (labeled as Single in Table 3), the performance of a fusion model for direct testing, the performance of a prompt model for direct testing and also show the ability of adjusting to shifted domain of LoRA.
Table 1: BroadUS-9.7K Dataset description.
Dataset Position Image Num Annotation
BUSI [1] Breast 780 Classification, Segmentation
BUSIS [22] Breast 562 Segmentation
UDIAT [20] Breast 163 Classification, Segmentation
BUS-BRA [5] Breast 1875 Classification, Segmentation
Fatty-Liver [2] Liver 550 Classification
kidneyUS [17] Kidney 534 Segmentation
DDTI [16] Thyroid 466 Segmentation
Fetal HC [8] Head 999 Segmentation
CAMUS [11] Cardiac 500 Segmentation
Appendix [14] Appendix 474 Classification
Table 2: Overall performance comparison on BoardUS-9.7K dataset.
Dataset BUS-BRA BUSIS CAMUS DDTI Fetal HC kidneyUS UDIAT Appendix BUS-BRA Fatty-Liver UDIAT Average
Task seg seg seg seg seg seg seg cls cls cls cls -
Single 80.44% 86.62% 89.80% 74.13% 85.36% 80.51% 75.02% 66.32% 73.07% 81.82% 69.70% 78.43%
Fusion 73.99% 87.39% 91.16% 61.15% 95.97% 82.29% 59.91% 53.68% 78.67% 90.91% 87.88% 78.46%
Prompt 75.59% 87.84% 92.05% 66.70% 96.69% 80.23% 61.00% 52.84% 84.69% 92.36% 88.79% 79.89%
The results of the first experiment are presented in Table 2. In general, the performance of single, pretrain, and prompt networks varied across individual datasets, with some exhibiting good performance while others performed poorly. However, the prompt method can consistently improved the average performance across all datasets. The average performance of the fusion model trained on the entire large dataset closely resembled that of the average performance of single model. These results demonstrate the efficacy of the fusion model framework and training strategy, as they are capable of handling diverse scenarios within the entire large dataset and achieving performance on par with specialized models. Remarkably, the prompt method provided an additional overall enhancement of 1.4%. These findings highlight the effectiveness of incorporating prompt information, which enhances both the capacity and capability of the network, resulting in improved performance across general tasks.
Table 3: Zero-shot domain shift performance comparison on BUSI dataset.
Dataset Task Scratch Single Fusion Prompt LoRA
BUSI seg 68.89% 60.85% 66.90% 66.91% 69.56%
BUSI cls 85.90% 80.80% 77.60% 78.40% 86.40%
The results of the second experiment are presented in Table 3. The table revealed that the zero-shot performance of the single model is unsatisfactory. The limited capacity and capability of a single model hinder its generalization performance. However, when comparing the prompt method to fusion, and fusion to single zero-shot results, there is an observable improvement in model performance. This indicates that models trained on a large number of datasets possess a certain level of generalization ability, and the effectiveness of prompt incorporation. It is noteworthy that even with just 50 epochs of fine-tuned training based on the Prompt architecture, the results surpass the performance achieved from training the model from scratch. This showcases the flexibility of our model and its ability to adapt to new datasets with minimal resources required for fine-tuning training.
Figure 3: t-SNE visualization.
Qualitative Analysis. We further employed t-SNE to visualize the distribution of representations. Specifically, we selected the breast datasets BUS-BRA, BUSIS, and UDIAT. The input to t-SNE was the features learned by the encoders of the single, fusion, and prompt models. As shown in Figure 3. There is clearly a noticeable domain shift in the single model result. In contrast, the fusion model results demonstrate a reduction in domain offset, with the datasets appearing more concentrated in the visualization. This indicates that the fusion model effectively mitigated the domain shift issue. Furthermore, the introduction of prompt messages resulted in nearly eliminating the domain offset. These findings align with the zero-shot results presented in Table 3. Finally, we visualized the segmentation results, as shown in Figure 4. The prompt model clearly outperformed the other models. This improvement can be attributed to the effective utilization of nature prompts and position prompts. By informing the model about the specific structure nature and its position, the prompt model achieved a deeper understanding of the segmentation task.
Figure 4: Some examples of segmentation result.Each column From left to right: original image, single, fusion, prompt, ground truth. The arrows show the details of how the edges are processed by different methods.
4 Conclusion
In this paper, we propose a universal model tailored for medical ultrasound imaging. Our model incorporates four distinct prompts, introducing information from various perspectives such as the input image nature, anatomical position, clinical task, and input image type. This comprehensive approach enhances the model’s learning capabilities. We conducted experiments using BroadUS-9.7K , a publicly available ultrasound dataset that we specifically collected. The results not only validate the effectiveness of our proposed framework but also demonstrate its potential for zero-shot testing. By expanding our analysis to include a wider range of datasets in the future, we aim to gain deeper insights into the effectiveness and adaptability of our proposed universal ultrasound model.